# Audio Array-Based 3D UAV Trajectory Estimation with LiDAR Pseudo-Labeling

### v1

@@ 1 Introduction
In recent years, Unmanned Aerial Vehicles (UAV) have become widely accessible, raising significant security concerns such as privacy violations, unauthorized surveillance, and risks in restricted airspace, including threats to civilian aviation. Their use in cross-border drug smuggling and military operations underscores the urgent need for effective UAV tracking technologies.
Despite the pressing need for UAV tracking systems, existing studies predominantly rely on visual information [1] [2] [3] [4] for 2D tracking or 3D detection from RADAR/LiDAR [5] [6] [7]. However, these approaches are limited by poor visibility, restricted field of view, and the high cost of RADAR/LiDAR for large-area surveillance and low-altitude smuggling UAV.
     The utilization of audio data, despite its accessibility and cost-effectiveness, remains largely unexplored [8, 9] in UAV trajectory estimation[10]. Given the affordability and ubiquity of microphones compared to LiDAR and cameras, there is a significant opportunity to develop innovative 3D tracking solutions that leverage audio information.

### v2

@@ 1 Introduction
In recent years, Unmanned Aerial Vehicles (UAV) [1] have become widely accessible, raising significant security concerns such as privacy violations [2], unauthorized surveillance [3], and risks in restricted airspace [4], including threats to civilian aviation. Their use in cross-border drug smuggling [5] and military operations underscores the urgent need for effective UAV tracking technologies [6].
Despite the pressing need for UAV tracking systems, existing studies predominantly rely on visual information [7] [8] [9] [10] for 2D tracking or 3D detection from RADAR/LiDAR [11] [12, 13] [14]. However, these approaches are limited by poor visibility, restricted field of view, and the high cost of RADAR/LiDAR for large-area surveillance and low-altitude smuggling UAV.
     The utilization of audio data, despite its accessibility and cost-effectiveness, remains largely unexplored [15, 16] in UAV trajectory estimation[17, 18]. Given the affordability and ubiquity of microphones compared to LiDAR and cameras [19], there is a significant opportunity to develop innovative 3D tracking solutions that leverage audio information.